1. SOLVE USING THE SUBSTITUTION METHOD.
x + y = 1
x - y = -3
2. Evaluate the determinant along row 1
a b
b a
3. Given matrix A, find the minor of the indicated element.
element: a 12
-2 -1 0
6 5 1
5 4 4
4. Determine the order of the matrix.
9 5 -3
-4 6 1



Answer by venugopalramana(3286)

1. SOLVE USING THE SUBSTITUTION METHOD.
x + y = 1.....Y=1-X...SUBSTITUTING
x - y = -3 .......X-(1-X)=-3....OR....X-1+X=-3....OR...2X=-3+1=-2...OR
X=-2/2=-1....SO....Y=1-(-1)=2
2. Evaluate the determinant along row 1
a b..
b a
=A*A*(-1)^(1+1)+B*B*(-1)(1+2)=A^2-B^2
WE TAKE EACH ELEMENT IN ROW 1,MULTIPLY WITH ITS COFACTOR AND ADD THEM UP.COFACTOR OF AN ELEMENT IS THE PRODUCT OF ITS MINOR AND (-1)^(ROW NUMBER + COLUMN NUMBER OF THE ELEMENT UNDER CONSIDERATION).MINOR OF AN ELEMENT IS THE DETERMINANT OR NUMBER OBTAINED BY DELETING THE ROW AND COLUMN CONTAINING THE ELEMENT UNDER CONSIDERATION ...HOPE THE ABOVE BREAK UP AND EXPLANATION IS CLEAR.
3. Given matrix A, find the minor of the indicated element.
element: a 12
-2 -1 0
6 5 1
5 4 4
WE HAVE TO FIND THE ELEMENT A12..THAT IS ELEMENT IN ROW 1 COLUMN 2...=-1
4. Determine the order of the matrix.
9 5 -3
-4 6 1
THERE ARE 2 ROWS AND 3 COLUMNS ...SO THE MATRIX IS OF THE ORDER (2,3)